OpenStudy (anonymous):

12 (x+4) = 3 (x+4)-9x please someone please help me who is good at explaining things. math isnt my strongest subject :((

OpenStudy (aroub):

hey...

OpenStudy (aroub):

well firs try distributing!

OpenStudy (aroub):

first*

OpenStudy (anonymous):

so 12 times 4. and 3 times 4 ???

OpenStudy (anonymous):

is that correct.

OpenStudy (aroub):

lets do the first one.. 12(x+4) . u multiply 12 by x and 12 by 4

OpenStudy (anonymous):

so now its 12 and 48 ???

OpenStudy (aroub):

no 12x and yes 24

OpenStudy (aroub):

48*

OpenStudy (aroub):

lol sryy

OpenStudy (aroub):

try the next one..

OpenStudy (anonymous):

okay. so 3x and 12.

OpenStudy (anonymous):

so putting that all together now do i have. 12x+48 = 3x+12-9x

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

helppp.

OpenStudy (aroub):

okay..yes ur right!! :)

OpenStudy (aroub):

now find x do u know how?

OpenStudy (anonymous):

OH MY. yayy : )

OpenStudy (aroub):

=)

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

umm im not sure what number to choose.

OpenStudy (aroub):

ok to do it in an easier put all that has a variable on a side and all that has a numbers on a side

OpenStudy (aroub):

like this: 12x-3x+9x=-48+12

OpenStudy (aroub):

now can u add them?

OpenStudy (anonymous):

just add up all of those numbers?

OpenStudy (aroub):

yes

OpenStudy (aroub):

try..

OpenStudy (anonymous):

oh i can do that.

OpenStudy (aroub):

ok quick,, b/c i have to go!

OpenStudy (anonymous):

-18.

OpenStudy (anonymous):

ah i hope that is right.

OpenStudy (aroub):

yes but u have to add the x

OpenStudy (aroub):

18x.

OpenStudy (anonymous):

after this im so giving you a medal btw.

OpenStudy (anonymous):

-18 !

OpenStudy (anonymous):

pppplllease tell me that is what x equals.

OpenStudy (aroub):

no its+18

OpenStudy (aroub):

lol u want the final answer?

OpenStudy (anonymous):

so 18x is the answer. yes?

OpenStudy (aroub):

yes

OpenStudy (anonymous):

thank you sooooooooo much :)

OpenStudy (anonymous):

you are a great teacher !

OpenStudy (aroub):

lol im not.. but thank you! =)

OpenStudy (aroub):

wait.. we still didnt finish!

OpenStudy (aroub):

18x=-36 u divide 18 by both sides so u get x= -36/18 x= -2

OpenStudy (anonymous):

i know you have to go. one LASt question where did that 36 come from.

OpenStudy (aroub):

sry yes.. -48+12 =-36 <--- this where -36 came from! :)